OpenStudy (moongazer):

3(sin 150degrees)(sin 90degrees) - 2(cos 210degrees)(cos 180degrees)

OpenStudy (anonymous):

3(0.5)(1)-2(-0.8)(-1) 1.5-1.6 = -0.1

OpenStudy (moongazer):

without using calculator :)

OpenStudy (moongazer):

I got 3/2 - sqrt of 3

OpenStudy (moongazer):

correct??

OpenStudy (anonymous):

3(1/2)(1)-2(-sqrt{3}/2)(-1) =3/2 minus .. hmm ..

OpenStudy (anonymous):

yeah i think so :)